A sample of students from an introductory psychology class were polled regarding the


number of hours they spent studying for the last exam. The data was used to make


inferences regarding the other students taking the course. There data are below:


8, 22.7, 14.5, 9, 9, 3.5, 8, 11, 7.5, 18,


20, 7.5, 9, 10.5,15, 19, 9, 8.5, 14,20


Compute a 99% confidence interval.

1
Expert's answer
2022-04-14T09:39:18-0400

The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{s}{\sqrt{n}},"

"CI=\bar{x}\pm z\cdot\cfrac{s}{\sqrt{n}},"

where:

  • "\bar{x}-" sample mean
  • z- the chosen z-value, for a 99% confidence interval z = 2.576
  • "s-" sample standard deviation
  • n =20 sample size.


The sample mean:

"\bar{x}=(8+ 22.7+ 14.5+ 9+ 9+ 3.5+ 8+ \\
+11+ 7.5+ 18+20+ 7.5+ 9+ 10.5+\\
+15+ 19+ 9+ 8.5+ 14+20)/ 20=12.19."


The sample variance:

"s_x^2=\cfrac{\sum(x_i-\mu)^2}{n-1},"

"s^2=((8-12.19)^2+(22.7-12.19)^2+(14.5-12.19)^2+...\\
+...(14-12.19)^2+(20-12.19)^2)/19=28.54."


The sample standard deviation:

"s=\sqrt{28.54}=5.34."


So,

"CI=12.19 \pm 2.576\cdot\cfrac{5.34}{\sqrt{20}}=\\=
12.19\pm3.08=(9.11, 15.27)."